Find the general solution of the given differential equation. y - 6y + 6y = Here y(t) =
Answer #1

el Ques yo-6y+64 = 0 characteristic ep of 82_68+6=0 =) = 6 + √36+24 2 = 6 + 512 = 6+2/3 = 3155 || (t) = x3t (a Cashist + G S